OpenStudy (daenio):

A bag contains 5 one-dollar tokens, 4 two-dollar tokens, 7 five-dollar tokens, and 6 ten-dollar tokens. Determine the expected value of a token, rounded to the nearest dollar. Question 5 options: 1 2 5 10

OpenStudy (daenio):

@Mertsj @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Expected value = Sum(probability*value) Expected value = (5/22)*1 + (4/22)*2 + (7/22)*5 + (6/22)*10 Expected value = 5/22 + 8/22 + 35/22 + 60/22 Expected value = 108/22 Expected value = 54/11 Expected value = 4.909 Expected value = 5 So the expected value is roughly $5

OpenStudy (daenio):

Thank you! :)

jimthompson5910 (jim_thompson5910):

you're welcome